Soybean meal is 16% protein and cormeal is 8% protein. How many pounds of each should be mixed together in order to get 320 pound mixture that is 13% protein?
How many pounds of cornemal should be in the mixture??__________
How many pounds of soybean meal should be in the mixture??__________
I am lost. can you please solve this for me!!!!

Answer by sarah1contreras(1)

Soybean meal is 16% protein. Cornmeal is 8% protein.
How many pounds of each should be mixed together in order to get 320
pound mixture that is 13% protein??
:
Let x = amt of soy
then the total is to be 320lb, therefore:
(320-x) = amt of corn
:
A typical mixture equation:
.16x + .08(320-x) = .13(320)
.16x + 25.6 - .08 = 41.6
.16x - .08x = 41.6 - 25.6
.08x = 16
x = 16/0.8
x = 200 lb of soy
then
320 - 200 = 120 lb of corn